OpenStudy (anonymous):

. An object's speed is 7.20 m/s, and its momentum is 360.0 kgm/s. What is the mass of the object?

OpenStudy (anonymous):

this is an easy one. momentum is equal to p=m*v we know that v=7.20 and momentum is 360.0, so plug those numbers into the equation: p=m*v 360=m*7.2 360/7.2=m m=50 kg